### v1

@@ 3.9.2 Predictive Factors for Model Success
Correlation analysis identified probability-accuracy calibration as the strongest individual predictor of overall model performance (r = 0.385, p = 0.005), followed by task-specific accuracy measures that were related to overall performance. Demographic variables, baseline VR performance, and subjective workload measures showed weak associations with model accuracy.
A predictive model incorporating demographic and performance variables explained only 13.7% of variance in individual model accuracy, indicating that readily observable participant characteristics provide limited predictive power for identifying suitable candidates for model-based adaptive control. The most informative predictor appears to be the probability calibration pattern itself, which can only be assessed through interaction with the system.
@@ 4.0.1 Key findings
ii) We found a clear trend towards participants in Experiment 2 (real-time deployment) judging the difficulty as Just Right more often than in the offline (non-adjusted) experiment 1. This trend was also true for the real-time adjusted dual task showing more Just Right judgments than a real-time adjusted Single Task. This is arguably the main contributor to the model performance drop as it was trained to distinguish the edge cases and was presented with a real-world dataset skewed towards Just Right targets.
@@
4.1 Adaptive Control Effectiveness Despite Prediction Limitations
@@
This apparent paradox of poor individual predictions but reasonable overall adaptation suggests that adaptive systems may be more robust to prediction errors than typically assumed. However, the model led to statistically harder difficulty levels in dual task conditions although with no cost on weighted performance. This indicates that participant control led to higher raw game performance on lower levels in dual task compared to the model which guided them to higher levels with lower raw game performance but matched weighted game performance.
@@ 4.4 Individual Differences and Model Effectiveness
The identification of probability calibration as the primary distinguishing factor between successful and unsuccessful participants provides crucial insights into the underlying mechanisms of physiological-based prediction. High-performing participants showed positive correlations between model class probabilities and prediction accuracy (r = 0.080), indicating that their physiological responses aligned systematically with the model’s learned patterns. Conversely, low-performing participants exhibited negative probability-accuracy correlations (r = -0.158), suggesting fundamental mismatches between their physiological signatures and the model’s training distribution.
This finding has implications for the deployment of adaptive VR systems. The probability calibration pattern suggests that brief screening interactions could identify suitable candidates for model-based control, potentially improving system effectiveness from the overall 36.7% to the 50.4% achieved by high-performing participants. Such user screening represents a viable pathway toward practical deployment, where probability calibration assessment becomes an integral component of adaptive training system design. We do not suggest excluding participants based on screening but rather identifying when the model should be recalibrated to result in a positive correlation.

### v2

@@ 3.9.2 Predictive Factors for Model Success
Correlation analysis identified probability-accuracy calibration as the strongest individual predictor of overall model performance (r=0.385,p=0.005). In the context of the zero-calibration deployment strategy, this metric served as a representation of the distributional alignment between an individual participant and the global training norms. Participants for whom the model performed best demonstrated positive correlations between model confidence and accuracy, indicating that their standardized physiological features aligned well with the population-level distributions used for scaling. Conversely, low-performing participants exhibited weak or negative correlations, suggesting a distributional shift where the global scaler failed to normalize their unique physiological baselines effectively.
In contrast to this physiological alignment metric, demographic variables (age, gender), baseline VR performance, and subjective workload measures showed only weak non-significant associations with model accuracy. A predictive model incorporating these readily observable characteristics explained only 13.7% of the variance in individual model accuracy. This indicates that traditional screening criteria provide limited predictive power for identifying suitable candidates for model-based adaptive control, as performance is primarily determined by the compatibility between the user’s physiology and the global normalization model.
@@ 4.0.1 Key findings
ii) We found a clear trend towards participants in Experiment 2 (real-time deployment) judging the difficulty as Just Right more often than in the offline (non-adjusted) experiment 1. This trend was confirmed by the significant reduction in polarized ratings (Too Easy and Too Difficult) in the real-time dataset compared to the offline baseline. As the model was specifically trained to detect these edge cases, the shift toward a predominantly Just Right distribution in the real-time experiment naturally degraded the classifier’s observed accuracy.
@@
ix) Our feature importance analysis revealed that the model’s decision-making was dominated by oculometric measures, suggesting that in active VR environments, immediate eye-movement dynamics provide a more robust signal for cognitive load adaptation than slower-acting physiological markers.
4.1 Adaptive Control Effectiveness
@@
This apparent paradox of poor individual predictions but reasonable overall adaptation suggests that adaptive systems may be more robust to prediction errors than typically assumed. In dual-task scenarios, we observed a ’conservatism bias’ in participant behavior. While the model pushed participants to significantly harder difficulty levels, this did not result in a degradation of difficulty-weighted performance or an increase in arithmetic errors compared to self-control. This suggests that when cognitive resources are divided, human users may underestimate their own capacity and self-regulate toward easier settings, whereas the objective model successfully identified that they could sustain higher challenge levels without performance costs.
@@
Finally, our time-series analysis demonstrates that the adaptive system successfully supported skill acquisition. Participants exhibited a continuous improvement in difficulty-weighted performance across the session, independent of the control type. Moreover, this performance improvement was accompanied by a significant reduction in subjective workload over time, particularly in dual-task conditions. This indicates that even with suboptimal individual prediction accuracy, the model-based adaptation fostered a flow-like state where users could learn and reduce their mental effort without being overwhelmed.
@@ 4.4 Individual Differences and Model Effectiveness
The identification of probability calibration as the primary predictor of success provides critical insight into the limitations of the global scaling strategy employed. High-performing participants showed positive correlations between model confidence and accuracy, indicating that their natural physiological baselines aligned with the global training distribution. Conversely, low-performing participants exhibited negative correlations, suggesting a distributional shift where the global scaler failed to normalize their unique physiological range effectively. This finding supports the argument that the probability-accuracy calibration observed here is a proxy for the physiological fit between the user and the global model. Consequently, this validates the necessity of a pre-task calibration phase - instead of screening users out, future systems must calculate individual baselines to align the user’s physiological distribution with the model’s decision boundaries prior to deployment.